OpenStudy (precal):

If g(x)=e^(2x)*f(x) and f(2)=-3, what is the equation of the normal line to the graph of g at x=2?

OpenStudy (precal):

\[g(x)=e ^{2x}*f(x)\]

OpenStudy (precal):

\[g'(x)=2e^{2x}*f(x)+ f ' (x)e^{2x}\]

OpenStudy (precal):

when x=2 what do I put for f ' (2)

OpenStudy (precal):

I might need to add a graph

OpenStudy (precal):

give me a moment to upload

OpenStudy (precal):

ok not sure that helps

OpenStudy (precal):

what are your thoughts?@hartnn

hartnn (hartnn):

that graph is of f'(x) so f'(2) is just the value of that curve at x=2

hartnn (hartnn):

f'(2) =4

OpenStudy (precal):

sorry can't see your prime did you put f ' (2)=4

hartnn (hartnn):

yes

OpenStudy (precal):

so the slope of the tangent line is -2e^4

OpenStudy (precal):

so the slope of the normal line is 1/(2e^4)

hartnn (hartnn):

correct

OpenStudy (precal):

ok \[y+3=\frac{ 1 }{ 2e^4 }(x-2)\]

hartnn (hartnn):

'y' value isn't -3

hartnn (hartnn):

for g(x) its g(2)

OpenStudy (precal):

oh so I have to calculate g(2), ok give me a moment

hartnn (hartnn):

yes normal line to graph of "g" so find a point on 'g' x=2, y =g(2)

OpenStudy (precal):

g(2)=-3e^4

OpenStudy (precal):

so the normal line is \[y-2=\frac{ 1 }{ 2e^4 }(x+3e^4)\]

OpenStudy (precal):

how does that look?

hartnn (hartnn):

you interchanged x1 and y1

hartnn (hartnn):

y- (-3e^4) = m (x-2)

OpenStudy (precal):

yes, I see that now....

OpenStudy (precal):

\[y+3e^4=\frac{ 1 }{ 2e^4 }(x-2)\]

OpenStudy (precal):

is that better?

hartnn (hartnn):

weird answer but out calculations seem to be correct...

OpenStudy (precal):

Can I ask you 4 more questions about that given f ' graph

hartnn (hartnn):

sure

OpenStudy (precal):

yes I have many weird answer problems

OpenStudy (precal):

1. On the interval [0,8], are they any values where f(x) is not differentiable? Give a reason for your answer.

OpenStudy (precal):

I want to say at x=2

OpenStudy (precal):

not sure why

hartnn (hartnn):

basically you cannot have a jump for a differentiable function

OpenStudy (precal):

I guess I can use the justification of a limit

hartnn (hartnn):

in other words, the slope should not change from + to - or - to + suddenly.

OpenStudy (precal):

ok I think I can justify it with the jump description or the limit definition.

OpenStudy (precal):

Next question, On what interval(s) is f " >0 f "(x)<0? Give reasons for your answers

OpenStudy (precal):

f "(x)>0 means f " (x) is positive correct? or that f (x) is concave up

hartnn (hartnn):

when f'(x) is increasing, f''(x) will be positive

OpenStudy (precal):

so f " (x) >0 on (0,2) and (6,8)

hartnn (hartnn):

yep

OpenStudy (precal):

and f " (x) < 0 on (4, -2)

OpenStudy (precal):

ok now I need to find two values f " (4) and f " (8) I don't know how to do this

hartnn (hartnn):

f " (x) < 0 on (2,5) and f''(x) =0 on (5,6)

OpenStudy (precal):

so am I just stating the slopes on those intervals

hartnn (hartnn):

note that f " (x) < 0 on (4, -2) was incorrect. i corrected it to f " (x) < 0 on (2,5)

hartnn (hartnn):

can you find the slope of line, given 2 points ?

OpenStudy (precal):

yes f " (4) = -2

hartnn (hartnn):

correct

OpenStudy (precal):

\[\frac{ f(5)-f(2) }{ 5-2 }=\frac{-2-4 }{ 3 }=\frac{ -6 }{ 3 }=-2\]

OpenStudy (precal):

what about f " (8)?

OpenStudy (precal):

I guess I should have written f ' (5) - f ' (2) for the numerator on top

hartnn (hartnn):

6,-2 8,2

OpenStudy (precal):

ok I can use those two points

hartnn (hartnn):

yes, thats correect f'(5) and f'(2)

OpenStudy (precal):

ok f " (8)=2

hartnn (hartnn):

yes

OpenStudy (precal):

Thanks you made it seem easy...

hartnn (hartnn):

it is easy :) welcome ^_^